OpenStudy (anonymous):

The formula to find the perimeter of a rectangle is: P = 2W + 2L. To solve for " W " in this formula, which steps should be followed? A.) Subtract 2L from both sides of the equation, then subtract 2 from both sides of the equation. B.) Subtract 2L from both sides of the equation, then divide both sides of the equation by 2. C.) Divide both sides of the equation by 2L, then subtract 2 from both sides of the equation. D.) Divide both sides of the equation by 2L + 2

OpenStudy (jackellyn):

P=2W+2L you want to get the W alone so Subtract 2L from both sides P-2L=2W and divide by 2 P-2L/2 =W P/2 - L = W

OpenStudy (anonymous):

so which one would that be ?

OpenStudy (anonymous):

So it's B

OpenStudy (jackellyn):

That is correct @rebecca1233